OpenStudy (anonymous):

x^2-2x=1 solve by completing the square

OpenStudy (anonymous):

x squared -2x=1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do I complete that square?

OpenStudy (anonymous):

Sorry but that I have no clue.

OpenStudy (anonymous):

Slight correction . x^2-2x+(-1)^2=1-1 x^2-2x+1=0 x=1